# VILA-U: a Unified Foundation Model Integrating Visual Understanding and Generation

### v2

@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated superior capabilities in various language tasks. Their appealing properties like instruction following, zero-shot generalization, and few-shot in-context learning motivate researchers to combine them with vision models to build visual language models (VLMs) for multi-modal tasks. Many efforts instructblip ; llava ; lin2023vila have been put into this field, achieving remarkable performance on visual language understanding benchmarks. In these works, visual inputs are projected onto LLMs’ semantic space through a vision foundation model like CLIP clip to bridge two modalities by including text-image alignment training objectives.
@@
In addition to visual understanding, another essential research direction in combining visual and language modalities is visual generation. There are two popular approaches for text-guided image generation. One approach employs diffusion models ldm , a powerful tool for various generation tasks. The other line of work converts visual content into discrete tokens through vector quantization (VQ) and then leveraging autoregressive transformers for high-quality and diverse generation VQGAN ; vit-vqgan ; lee2022autoregressive .
@@
Witnessing the rapid advancements in both visual understanding and generation, an emerging trend is to unify these techniques into a single multi-modal framework. There are two main approaches to achieving such unification. Many VLMs jin2023unified ; li2024mgm ; Emu ; Emu2 maintain an understanding-oriented framework and offload the generation task to an external diffusion model. This disjoint approach adds complexity to infrastructure design. Available large-scale foundation model training pipelines and deployment systems have already been highly optimized for language modeling with next-token prediction. Designing a new stack to support diffusion models would incur significant engineering costs. To circumvent such costs, it is desirable to design a single end-to-end autoregressive framework for both image understanding and generation. There is a trend in VLMs lwm ; CM3Leon that adopt VQ encoders to convert visual inputs into discrete tokens and treat them in the same next-token prediction manner as language data. However, replacing continuous tokens with VQ tokens in VLMs usually results in a severe performance drop in downstream visual perception tasks. Other works Unified-io2 ; chameleon have to make various architectural modifications and conduct multi-modal training from scratch, which is computationally expensive.
@@
In this work, we present VILA-U, an end-to-end autoregressive framework with a unified next-token prediction objective for both visual and text inputs that can achieve competitive performance on both visual language understanding and generation tasks, without the help of external components like diffusion models. We identify two critical principles to unify vision and language modalities effectively and efficiently. (1) Existing end-to-end autoregressive VLMs cannot achieve competitive visual understanding performance because the discrete VQ tokens are trained solely on image reconstruction loss and are not aligned with textual inputs. Therefore, it is crucial to introduce text alignment during VQ vision tower pretraining to enhance perception capabilities. (2) Autoregressive image generation can attain similar quality as diffusion models if trained on a high-quality data corpus with sufficient size. Guided by these insights, VILA-U features a unified foundation vision tower that converts visual inputs into discrete tokens through vector quantization and aligns these tokens with textual inputs using contrastive learning. The multi-modal training of VILA-U takes advantage of a unified next-token prediction objective for both visual and textual tokens on a small-size high-quality image-text corpus.
@@ 2 Related Work
Large Language Models (LLMs). LLMs based on pre-trained large-scale transformers NIPS2017Transformer has drastically revolutionized natural language processing field. Featuring gigantic model size and pre-training data corpus, LLM has achieved remarkable performance on various linguistic tasks. The development of open-source LLMs such as LLaMA llama , Mixtral mixtral and Vicuna vicuna has furthered nourished research on how to adopt LLM for complex language tasks. Besides excellent zero-shot generalizability to diverse domains, LLM is commonly finetuned on custom datasets for better performance on specific tasks. Instruction tuning ChatGPT ; chung2024scaling ; ouyang2022training also stands as a key step for better outputs in applying LLMs. In this work, we adopt the LLaMA-2-7Bllama model as our basic LLM.
Visual Language Models (VLMs). Combining computer vision and natural language processing gives rise to VLM in this LLM era. In VLMs, researchers leverage vision foundation models such as CLIP clip , BLIP blip and CoCa CoCa to extract visual features, align with texts, and feed them into LLM to achieve the cross-modality understanding between texts and visual content. Building upon such progress, many VLMs alayrac2022flamingo ; blip2 ; llava ; lin2023vila have been designed and trained on extensive vision-language data to achieve remarkable performance on visual understanding and reasoning tasks. VLMs are divided into two types. (1) BLIP-style VLMs openflamingo ; alayrac2022flamingo ; li2022blip ; li2023blip ; Dai2023InstructBLIP ; hong2023cogagent utilizes cross attention mechanism to fuse language and visual information and optionally apply perceivers jaegle2021perceiver to downsample visual tokens. (2) LLaVA-style VLMs liu2023llava ; driess2023palm ; chen2023pali ; fuyu ; zhu2023minigpt ; ye2023mplug ; qwen ; aiello2023jointly ; chen2023internvl ; liu2023improved ; lin2023vila ; zhang2023internlm converts visual inputs to tokens (patches) and pass them through ViTs. The output of ViTs undergoes MLP layers and gets aligned to the language space. In this work, we aim to develop a VLM with visual understanding capacities comparable to prior works, while also possessing the new capacity of visual generation.
Unified Visual Language Models. Numerous efforts have been made to develop unified visual language models capable of generating both text and visual content, including images and videos. There are two mainstream methods to generate visual content in VLMs. Many works Emu ; Emu2 ; jin2023unified ; video-lavit ; li2024mgm combine VLMs with diffusion models like Stable Diffusion ldm for high-quality image generation. Other works lwm ; CM3Leon ; Unified-io2 ; chameleon ; xie2024show adopt VQGAN-based vision encoders to convert visual inputs into discrete tokens and make LLMs learn to predict them. In this work, we design our framework based on the autoregressive next-token prediction method for visual generation and make our VLM learn to generate visual content effectively and efficiently.
@@ 3 Methods
This work proposes a multi-modal framework that aims to unify visual and language modalities efficiently and effectively. The key components enabling such unification are a unified foundation vision tower that converts visual inputs into discrete tokens aligned with text, and a unified multi-modal generative training procedure. An overview of the main multi-modal training and inference process within our framework is depicted in Figure 1.
@@ 3.1 Unified Foundation Vision Tower
A straightforward combination of contrastive and reconstruction loss cannot converge. This is because alignment and reconstruction tasks require high-level semantic and low-level appearance features, respectively. Training the entire vision tower from scratch with both objectives could induce conflicting goals. In practice, we observe that training the vector-quantized vision tower from scratch with both image reconstruction and contrastive loss results in a mere 5% Top-1 accuracy for zero-shot image classification on ImageNet deng2009imagenet after several epochs of training.
To address this issue, we experiment with different training recipes and find the following solution to be most effective. Instead of learning both objectives simultaneously, our training recipe suggests first equipping the model with text-image alignment ability and then learning reconstruction while maintaining alignment ability. We initialize the vision encoder and text encoder with pretrained weights from the CLIP model to ensure good text-image alignment. Next, we freeze the text encoder and keep all vision components trainable using both contrastive and reconstruction loss. The contrastive loss maintains alignment ability, while the reconstruction loss develops reconstruction ability. This training approach converges quickly and yields strong performance. The pre-trained CLIP weights contain learned high-level priors, which are difficult and computationally expensive to learn from scratch. Initializing with these weights enables the binding of low-level and high-level features much faster and more tractably for the vision encoder. With this training recipe, we can efficiently train a vision tower that exhibits both good text alignment and image reconstruction abilities. We use weighted sum to combine the text-image contrastive loss and VQ-based image reconstruction loss:
@@
Discussion: Failed Training Recipes.
We experiment with numerous training recipes and find none to be as effective as our final approach. We list four alternative recipes and discuss their shortcomings compared to our final recipe: (1) Load pre-trained CLIP weights into the text encoder only; (2) Load pre-trained RQ-VAE weights for the vision encoder and decoder while training other parts from scratch; (3) Freeze the vision encoder; (4) Make the text encoder trainable.
@@
Our visual features are discretely quantized, so their representation ability heavily depends on the code size used in our quantizer. Since we hope they contain both high-level and low-level features, we need more capacities in their vector feature space, making a larger code size necessary for good performance in downstream tasks. However, too many codes for each image will result in too many tokens for LLM to produce in the visual generation process, incurring much latency. So in an attempt to increase the vector feature capacity and meanwhile maintain a reasonable number of tokens for LLM, we adopt a residual vector quantization method following RQ-VAE lee2022autoregressive to discretize a vector $\mathbf{z}$ as $D$ discrete codes:
@@ 3.2 Unified Multi-modal Generative Pre-training
where $T$ is the length of the multi-modal sequence and $i$ only counts when the text token appears at position $i$. For visual tokens, residual quantization introduces a depth-stacked structure of codes at each visual position $j$. To address this, we leverage the depth transformer introduced in RQ-VAE lee2022autoregressive . Specifically, given the code embedding $h_{j}$ generated by the LLM for visual tokens at position $j$, the depth transformer autoregressively predicts D residual tokens ($k_{j1}$, …, $k_{jD}$). During training, the input of the depth transformer $v_{jd}$ at depth d is defined as the sum of the code embeddings of up to depth $d-1$ for $d>1$ such that
@@
and $v_{j1}=h_{j}$. Thus, the depth transformer predicts the next code for a finer estimation of the feature $\boldsymbol{\hat{z}}_{j}$ based on the previous estimations up to $d-1$. Then the negative log-likelihood loss for visual tokens is
@@ 4 Experiments
In this section, we introduce comprehensive experiments to evaluate the performance of our method on various visual understanding and generation tasks. Firstly, we outline our experimental setup, including the model architecture, training datasets, and evaluation benchmarks. Subsequently, we evaluate the performance of our unified foundation vision tower. Then, we compare our method with other popular VLMs on various visual understanding and generation benchmarks. Finally, we give some qualitative results.
@@ 4.1 Experimental Setup
In our experiments, we employ LLaMA-2-7B touvron2023llama as our base language model. For the vision tower, we choose SigLIP-Large-patch16-256 / SigLIP-SO400M-patch14-384 zhai2023sigmoid as our vision encoder architecture, and adopt the residual quantizer, depth transformer as well as the decoder architecture from RQ-VAE lee2022autoregressive . The quantizer codebook size is 16384. All images and videos are resized to a resolution of $256\times 256$ / $384\times 384$, with each image or video frame converted into a $16\times 16\times 4$ / $27\times 27\times 16$ code with the residual depth $D=4$ / $D=16$. We train our vision tower on COYO-700M coyo-700m and evaluate it for zero-shot classification and reconstruction performance on ImageNet imagenet . For visual understanding, we leverage 1M [image, text] data from ShareGPT4V sharegpt4v , 6M interleaved text and image data from MMC4 mmc4 . For visual generation, we incorporate 15M high-quality [text, image] data curated from our internal dataset and 1M [text, video] data from OpenVid nan2024openvid datasets. Classifier-free guidance cfg is employed for visual generation with a CFG value of 3.
For examining visual understanding ability, we evaluate our model on the widely adopted zero-shot image-based visual-language benchmarks including VQA-v2 vqa_v2 , GQA gqa , TextVQA textvqa , POPE POPE , MME mme , SEED seed , MM-Vet mm-vet and video-based visual-language benchmarks including ActivityNet activitynet , MSVD chen2011collecting , MSRVTT MSVD , TGIF tgif .
To evaluate the visual generation capability, we use MJHQ-30K playgroundv2.5 and GenAI-Bench GenAI-Bench as our benchmarks. The former adopts the FID between generated images and 30K high-quality images to reflect the overall capability of image generation. The latter is a challenging image-to-text generation benchmark that reflects the comprehensive generative abilities of visual generation models. This benchmark is divided into two categories of prompts: basic skills, which include attribute, scene, and relation understanding in text inputs, and advanced skills, which encompass counting, differentiation, comparison, and logical relation understanding in text inputs.
@@ 4.2 Unified Foundation Vision Tower
We present the commonly used metrics reconstruction FID (rFID) and Top-1 accuracy for zero-shot image classification on ImageNet to measure the reconstruction and text alignment capabilities of the unified foundation vision tower in Table 1. Our model achieves significantly better reconstruction results than VQ-GAN. Our rFID is slightly inferior to that of RQ-VAE when using the same code shape. This is expected as the introduction of contrastive loss during training, aimed at enhancing image understanding, led to a decrease in reconstruction quality. For the text alignment capability, our unified vision tower achieves a Top-1 accuracy of 73.3 / 78.0 under 256 / 384 resolution. This demonstrates the exceptional text alignment capability of our unified vision tower. However, it is worth noting that both the rFID and Top-1 accuracy of the vision tower only serves as a medium indicator instead of directly linear correlated to the final performance of our whole multi-modal framework. The performance on visual understanding and generation tasks presented in the following sections is more important.
@@
VQ-GAN esser2021taming 256 $\times$ 256 16 $\times$ 16 4.98
RQ-VAE lee2022autoregressive 256 $\times$ 256 8 $\times$ 8 $\times$ 4 3.20
RQ-VAE lee2022autoregressive 256 $\times$ 256 16 $\times$ 16 $\times$ 4 1.30

### v3

@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated superior capabilities in various language tasks. Their appealing properties like instruction following, zero-shot generalization, and few-shot in-context learning motivate researchers to combine them with vision models to build visual language models (VLMs) for multi-modal tasks. Many efforts (Dai et al., 2024; Liu et al., 2024b; Lin et al., 2023) in this field have achieved remarkable performance on visual language understanding. In these works, visual inputs are projected onto LLMs’ semantic space through a vision model like CLIP (Radford et al., 2021) to bridge two modalities by including text-image alignment objectives.
@@
In addition to visual understanding, another essential research direction in combining visual and language modalities is visual generation. There are two popular approaches for text-guided image generation. One approach employs diffusion models (Rombach et al., 2022a), a powerful tool for various generation tasks. The other line of work converts visual content into discrete tokens through vector quantization (VQ) and then leveraging autoregressive transformers for high-quality and diverse generation (Esser et al., 2021; Yu et al., 2021; Lee et al., 2022; Tian et al., 2024b; Sun et al., 2024).
@@
Witnessing the rapid advancements in both visual understanding and generation, an emerging trend is to unify these techniques into a single multi-modal framework. Prior to VILA-U, there are two main approaches to achieving such unification: (1) One approach (Liu et al., 2024a; Yu et al., 2023a; Xie et al., 2024) utilizes a VQGAN-based (Esser et al., 2021) tokenizer to convert visual inputs into discrete tokens and leverages an autoregressive model for both understanding and generation. However, (Xie et al., 2024) has shown that visual tokens from VQGAN-based encoder lack semantic information and usually results in a severe performance drop in downstream visual understanding tasks. (2) Another approach (Zhan et al., 2024; Ge et al., 2023b; Jin et al., 2023) utilizes a codebook to quantize features produced by a pre-trained vision model like CLIP. Since CLIP features encode rich semantic information, these approaches generally achieve significantly better performance on understanding tasks. However, these tokenizers lack decoding capability, requiring an external visual generation model, such as a diffusion model, to use the generated visual tokens as conditions for producing visual outputs. This approach adds complexity to infrastructure design. Available large-scale foundation model training pipelines and deployment systems have already been highly optimized for language modeling with next-token prediction. Designing and maintaining an additional stack to support diffusion models would incur significant engineering costs.
@@
In this work, we present VILA-U, an end-to-end autoregressive framework with a unified next-token prediction objective for both visual and text inputs that can achieve competitive performance on both visual language understanding and generation tasks, without the help of external components like diffusion models. We identify two critical principles to unify vision and language modalities: (1) Existing unified end-to-end autoregressive VLMs cannot achieve competitive visual understanding performance because the discrete VQGAN tokens are trained solely on image reconstruction loss and are not aligned with textual inputs. Therefore, it is crucial to introduce text alignment during VQ vision tower pretraining to enhance perception capabilities. (2) Autoregressive image generation can attain similar quality as diffusion models if trained on high-quality data with sufficient size. Guided by these insights, VILA-U features a unified foundation vision tower that converts visual inputs into discrete tokens through vector quantization and aligns these tokens with textual inputs using contrastive learning. The multi-modal training of VILA-U takes advantage of a unified next-token prediction objective for both visual and textual tokens on a small-size high-quality image-text corpus.
@@ 2 Related Work
Large Language Models (LLMs). LLMs based on pre-trained large-scale transformers (Vaswani et al., 2017) has drastically revolutionized natural language processing field. Featuring gigantic model size and pre-training data corpus, LLM has achieved remarkable performance on various linguistic tasks. The development of open-source LLMs such as LLaMA (Touvron et al., 2023a), Mixtral (Jiang et al., 2024) and Vicuna (Chiang et al., 2023) has furthered nourished research on how to adopt LLM for complex language tasks. Besides excellent zero-shot generalizability to diverse domains, LLM is commonly finetuned on custom datasets for better performance on specific tasks. Instruction tuning (OpenAI, 2023; Chung et al., 2024; Ouyang et al., 2022) also stands as a key step for better outputs in applying LLMs. In this work, we adopt the LLaMA-2-7B (Touvron et al., 2023a) model as our basic LLM.
Visual Language Models (VLMs). Combining computer vision and natural language processing gives rise to VLM in this LLM era. In VLMs, researchers leverage vision foundation models such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022) and CoCa (Yu et al., 2022) to extract visual features, align with texts, and feed them into LLM to achieve the cross-modality understanding between texts and visual content. Building upon such progress, many VLMs (Alayrac et al., 2022; Li et al., 2023b; Liu et al., 2024b; Lin et al., 2023; Luo et al., 2024; Tian et al., 2024a) have been designed and trained on extensive vision-language data to achieve remarkable performance on visual understanding and reasoning tasks. In this work, we aim to develop a VLM with visual understanding capacities comparable to prior works, while also possessing the new capacity of visual generation.
Unified Visual Language Models. Numerous efforts have been made to develop unified visual language models capable of generating both text and visual content, including images and videos. There are two mainstream methods to generate visual content in VLMs. Many works (Sun et al., 2023b; a; Jin et al., 2023; Ge et al., 2023b; Li et al., 2023c; Ge et al., 2024; Jin et al., 2024; Ge et al., 2023a) combine VLMs with diffusion models like Stable Diffusion (Rombach et al., 2022a) for high-quality image generation. Other works (Liu et al., 2024a; Yu et al., 2023a; Lu et al., 2023; Team, 2024; Xie et al., 2024) adopt VQGAN-based vision encoders to convert visual inputs into discrete tokens and make LLMs learn to predict them. For more details on the distinction between our method and other unified visual language models, please refer to Appendix A.
@@ 3 Methods
This work proposes a multi-modal framework that aims to unify visual and language modalities effectively. The key components enabling such unification are a unified foundation vision tower that converts visual inputs into discrete tokens aligned with text, and a unified multi-modal generative training procedure. An overview of the main multi-modal training and inference process within our framework is depicted in Figure 1.
@@ 3.1 Unified Foundation Vision Tower
Training the unified vision tower with two objectives from scratch would be difficult, because alignment and reconstruction tasks require high-level semantic and low-level appearance features, respectively. Training the entire vision tower from scratch with both objectives could induce conflicting goals. In practice, we observe that training the vector-quantized vision tower from scratch with both image reconstruction and contrastive loss results in a mere 5% Top-1 accuracy for zero-shot image classification on ImageNet (Deng et al., 2009a) after several epochs of training.
To address this issue, we experiment with different training recipes (failed recipes are listed in Appendix C) and find the following solution to be most effective. Instead of learning both objectives simultaneously, we suggest first equipping the model with text-image alignment ability and then learning reconstruction while maintaining alignment ability. We initialize the vision encoder and text encoder with pretrained weights from the CLIP model to ensure good text-image alignment. Next, we freeze the text encoder and keep all vision components trainable using both contrastive and reconstruction loss. The contrastive loss maintains alignment ability, while the reconstruction loss develops reconstruction ability. This approach converges quickly and yields strong performance. The pre-trained CLIP weights contain learned high-level priors, which are difficult and computationally expensive to learn from scratch. Initializing with these weights enables the binding of low-level and high-level features much faster and more tractably for the vision encoder. With this recipe, we can train a vision tower that exhibits both good text alignment and image reconstruction abilities. We use weighted sum to combine the text-image contrastive loss and VQ-based image reconstruction loss:
@@
Our visual features are discretely quantized, so their representation ability heavily depends on the code size used in our quantizer. Since we hope they contain both high-level and low-level features, we need more capacities in their vector feature space, making a larger code size necessary for good performance in downstream tasks. However, too many codes for each image will result in too many tokens for LLM to produce in the visual generation process, incurring much latency. So in an attempt to increase the vector feature capacity and meanwhile maintain a reasonable number of tokens for LLM, we adopt a residual vector quantization method following RQ-VAE (Lee et al., 2022) to discretize a vector $\mathbf{z}$ as $D$ discrete codes:
@@ 3.2 Unified Multi-modal Generative Pre-training
where $T$ is the length of the multi-modal sequence and $i$ only counts when the text token appears at position $i$. For visual tokens, residual quantization introduces a depth-stacked structure of codes at each visual position $j$. To address this, we leverage the depth transformer introduced in RQ-VAE (Lee et al., 2022). Specifically, given the code embedding $h_{j}$ generated by the LLM for visual tokens at position $j$, the depth transformer autoregressively predicts D residual tokens ($k_{j1}$, …, $k_{jD}$). During training, the input of the depth transformer $v_{jd}$ at depth d is defined as the sum of the code embeddings of up to depth $d-1$ for $d>1$ such that
@@
and $v_{j1}=h_{j}$. Thus, the depth transformer predicts the next code for a finer estimation of the feature $\bm{\hat{z}}_{j}$ based on the previous estimations up to $d-1$. Then the negative log-likelihood loss for visual tokens is
@@ 4 Experiments
In this section, we introduce comprehensive experiments to evaluate our method on various visual understanding and generation tasks. Firstly, we outline our experimental setup, including the model architecture, training datasets, and evaluation benchmarks. Subsequently, we evaluate the performance of our unified foundation vision tower. Then, we compare our method with other popular VLMs on various visual understanding and generation benchmarks. Finally, we give some qualitative results.
@@ 4.1 Experimental Setup
In our experiments, we employ LLaMA-2-7B (Touvron et al., 2023b) as our base language model. For the vision tower, we choose SigLIP-Large-patch16-256 / SigLIP-SO400M-patch14-384 (Zhai et al., 2023) as our vision encoder architecture, and adopt the residual quantizer, depth transformer as well as the decoder architecture from RQ-VAE (Lee et al., 2022). The quantizer codebook size is 16384. All images and videos are resized to a resolution of $256\times 256$ / $384\times 384$, with each image or video frame converted into a $16\times 16\times 4$ / $27\times 27\times 16$ code with the residual depth $D=4$ / $D=16$. We train our vision tower on COYO-700M (Byeon et al., 2022) and evaluate it for zero-shot classification and reconstruction performance on ImageNet (Deng et al., 2009b). For visual understanding, we leverage 1M [image, text] data from ShareGPT4V (Chen et al., 2023), 6M interleaved text and image data from MMC4 (Zhu et al., 2024). For visual generation, we incorporate 15M high-quality [text, image] data curated from our internal dataset and 1M [text, video] data from OpenVid (Nan et al., 2024) datasets. Classifier-free guidance (Ho & Salimans, 2022) is employed for visual generation with a CFG value of 3.
For examining visual understanding ability, we evaluate our model on the widely adopted zero-shot image-based visual-language benchmarks including VQAv2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), TextVQA (Singh et al., 2019), POPE (Li et al., 2023d), MME (Fu et al., 2024), SEED (Li et al., 2023a), MM-Vet (Yu et al., 2023b) and video-based visual-language benchmarks including ActivityNet (Caba Heilbron et al., 2015), MSVD (Chen & Dolan, 2011), MSRVTT (Xu et al., 2017), TGIF (Li et al., 2016).
To evaluate the visual generation capability, we use MJHQ-30K (Li et al., 2024) and GenAI-Bench (Lin et al., 2024) for image generation and VBench (Huang et al., 2024) for video generation. MJHQ-30K adopts the FID between generated images and 30K high-quality images to reflect the overall capability of image generation. GenAI-Bench is a challenging image-to-text generation benchmark that reflects the comprehensive generative abilities of image generation models. Vbench is a comprehensive benchmark suite for video generative models that decomposes the generation quality into multiple well-defined dimensions to facilitate fine-grained and objective evaluation.
@@ 4.2 Unified Foundation Vision Tower
We present the commonly used metrics reconstruction FID (rFID) and Top-1 accuracy for zero-shot image classification on ImageNet to measure the reconstruction and text alignment capabilities of the unified foundation vision tower in Table 1. Please refer to the Appendix B.1 for the qualitative reconstruction results. Our model achieves significantly better reconstruction results than VQ-GAN. Our rFID is slightly inferior to that of RQ-VAE when using the same code shape. This is expected as the introduction of contrastive loss during training, aimed at enhancing image understanding, led to a decrease in reconstruction quality. For the text alignment capability, our unified vision tower achieves a Top-1 accuracy of 73.3 / 78.0 under 256 / 384 resolution. This demonstrates the exceptional text alignment capability of our unified vision tower. However, it is worth noting that both the rFID and Top-1 accuracy of the vision tower only serves as a medium indicator. As the unified vision tower is an integral component of the entire autoregressive model, we believe that its performance on downstream tasks, such as visual understanding and generation, holds greater significance.
@@
VQ-GAN 256 $\times$ 256 16 $\times$ 16 4.98
RQ-VAE 256 $\times$ 256 8 $\times$ 8 $\times$ 4 3.20
RQ-VAE 256 $\times$ 256 16 $\times$ 16 $\times$ 4 1.30